# DINeMo: Learning Neural Mesh Models with no 3D Annotations

### v1

@@ Abstract
Category-level 3D/6D pose estimation is a crucial step towards comprehensive 3D scene understanding, which would enable a broad range of applications in robotics and embodied AI. Recent works explored neural mesh models that approach a range of 2D and 3D tasks from an analysis-by-synthesis perspective. Despite the largely enhanced robustness to partial occlusion and domain shifts, these methods depended heavily on 3D annotations for part-contrastive learning, which confines them to a narrow set of categories and hinders efficient scaling. In this work, we present DINeMo, a novel neural mesh model that is trained with no 3D annotations by leveraging pseudo-correspondence obtained from large visual foundation models. We adopt a bidirectional pseudo-correspondence generation method, which produce pseudo correspondence utilize both local appearance features and global context information. Experimental results on car datasets demonstrate that our DINeMo outperforms previous zero- and few-shot 3D pose estimation by a wide margin, narrowing the gap with fully-supervised methods by 67.3%. Our DINeMo also scales effectively and efficiently when incorporating more unlabeled images during training, which demonstrate the advantages over supervised learning methods that rely on 3D annotations. Our project page is available here.
@@ 1 Introduction
In this work, we present DINeMo, a novel neural mesh model that is trained without 3D annotations by leveraging large pretrained visual foundation models, such as DINOv2 [20]. Rather than relying on groundtruth keypoint locations as in previous works [13, 9], we train DINeMo with part-contrastive loss using pseudo-correspondence obtained by matching neural features from SD-DINO [35]. However, raw pseudo-correspondence from SD-DINO can be quite noisy, e.g., keypoints are often mismatched between left and right. We argue that keypoint correspondence matching should consider both local information, i.e., per patch feature similarities, and global context information, i.e., 3D orientation of the object. Based on this motivation, we propose a novel bidirectional pseudo-correspondence generation, which consists of two steps: (i) matching a global pose label from raw keypoint correspondences, and (ii) refine local keypoint correspondences based on the predicted global pose label. Lastly we extend standard analysis-by-synthesis inference with Grounded-SAM[23] masks, achieving enhanced occlusion robustness.
Extensive results on car class demonstrate that our DINeMo outperforms previous zero- and few-shot 3D pose estimation methods by a wide margin on both in-distribution and partial occlusion testing data, significantly narrowing the gap from fully-supervised methods by 67.3%. Our approach also outperforms all previous methods on SPair71k [16] for semantic correspondence.
As DINeMo can be trained on object images without 3D annotations, we further study the scaling properties of our approach by involving abundant data from public image datasets, such as Stanford Cars [12]. Experimental results demonstrate that our DINeMo scales effectively efficiently by involving more unlabeld images for training. This demonstrates the advantages of DINeMo over previous fully-supervised approaches [13, 9], which struggle to scale up due to the reliance on 3D annotations.
In summary, our main contributions are as follows: (1) We present DINeMo, a novel neural mesh model trained with pseudo-labels from our bidirectional pseudo-correspondence generation method. (2) Experimental results demonstrate that DINeMo outperforms previous zero- and few-shot methods by a wide margin, largely narrowing the gap with fully-supervised methods. (3) Our DINeMo scales effectively and efficiently with more unlabeled images during training.
@@ 2 Related Work
Multiple strategies have been developed for category-level 3D pose estimation. Traditional methods treat this as a classification problem, assigning discrete pose labels to objects [28, 17]. Another line of work follows a two-stage, keypoint-driven pipeline [39], where semantic keypoints are detected and used to estimate 3D pose via Perspective-n-Point algorithm. More recent advancements shifted toward render-and-compare paradigms [30, 13, 9, 33], transitioning from generating pixel values to semantic feature representations. These methods frame pose estimation as an optimization problem that minimizes the discrepancy between features extracted from the input image and ones rendered from a posed 3D mesh.
Keypoint correspondence. Several approaches have been developed to establish correspondences between 2D image keypoints and 3D mesh vertices—a fundamental component in tasks such as shape reconstruction and 3D pose estimation. At the category level, however, most existing methods focused on deformable object classes, particularly humans [4, 22] and animals [26]. This focus is largely attributed to the availability of large-scale datasets with dense image-to-template annotations, such as DensePose-COCO [4] for humans and DensePose-LVIS [19] for animals. Notably, recent work has demonstrated the feasibility of learning dense correspondences without explicit keypoint supervision [26], particularly in the animal domain. In this work, we extend this line of research to rigid object categories, which have received comparatively less attention.
Render-and-compare approaches estimate 3D poses by minimizing the reconstruction error between feature representations projected from a 3D object and those extracted from the input image. This strategy can be interpreted as a form of approximate analysis-by-synthesis [3], which contrasts with purely discriminative methods and has demonstrated increased robustness to out-of-distribution scenarios. Particularly, such approaches have shown effectiveness in handling partial occlusions, proving beneficial in both object classification [11] and 3D pose estimation tasks [30, 8].
@@ 3 Methods
In this section we introduce our DINeMo, as illustrated in Fig. 1. We start by reviewing the neural mesh models in Sec. 3.1. Then we introduce our bidirectional pseudo-correspondence generation in Sec. 3.2. Lastly we present our inference method in Sec. 3.3, which
@@ 3.1 Preliminaries: Neural Mesh Models
where $f_{\Phi}$ is the network backbone parameterized by $\Phi$, $\mathcal{FG}$ and $\mathcal{BG}$ are set of foreground and background positions, and $C_{b}$ is the background feature. The network parameters $\Phi$ and learnable feature $\{\mathcal{C},b\}$ are optimized with a part-contrastive loss. During inference, we minimize the negative log-likelihood w.r.t. pose parameters $m$.
@@ 3.2 Bidirectional Pseudo-Correspondence
Previous works train neural mesh models on keypoint correspondences obtained from 3D annotations, such as object poses or human poses. In this work, we propose to leverage pseudo-correspondence from visual foundation models, without relying on 3D annotations that are hard to obtain.
@@ 3.3 Inference
To find foreground and background regions, previous works [13, 9, 30] introduce an one-hot map $\mathbf{Z}$ indicating if each 2D location is visible or not based on feature activations. Empirically we find that the one-hot map often involves a significant portion of the background, leading to degraded performance on partial occlusion data. Given the recent advancements of segmentation methods, we first predict the $\mathcal{Z}$ with a SAM2 [21] model and then optimizes the negative log-likelihood loss while fixing $\mathbf{Z}$.
@@ 4.2 Scaling Properties
As our DINeMo is trained on object images without 3D annotations, we study the scaling properties w.r.t. different training data sizes. Specifically, we trained a variety of DINeMo models on different number of unlabeled images, ranging from 2048 images (comparable to the training set size in PASCAL3D+ [32]) up to 15,000 images (as in Stanford Cars dataset [12]). As shown from the results in Fig. 4, DINeMo scales well with more unlabeled images used during training, i.e., the pose accuracy at $\pi/6$ increases from 93.1% to 93.7% and the per point PCK increases from 62.3 to 64.2. This highlights the advantages over previous supervised learning methods — DINeMo does not require 3D annotations that are hard to obtain, and scales effectively and efficiently by involving more unlabeled images for training, which are abundant on the Internet [25].
@@ 5 Conclusions
In this work we present DINeMo, a novel neural mesh model that is trained with no 3D annotations-only from pseudo-correspondence obtained from large visual foundation models. We propose a novel bidirectional pseudo-correspondence generation method that can effectively utilize both local appearance features and global context information to produce more 3D-consistent pseudo-correspondence. Experimental results on car datasets demonstrate that our DINeMo outperforms previous zero- and few-shot 3D pose estimation methods by a wide margin, narrowing the gap with fully-supervised methods by 67.3%. By incorporating more unlabeled images during training, our DINeMo also scales effectively and efficiently, demonstrating the advantages over fully supervised methods tha rely on scarce 3D annotations.

### v2

@@ Abstract
Category-level 3D/6D pose estimation is a crucial step towards comprehensive 3D scene understanding, which would enable a broad range of applications in robotics and embodied AI. Recent works explored neural mesh models that approach a range of 2D and 3D tasks from an analysis-by-synthesis perspective. Despite the largely enhanced robustness to partial occlusion and domain shifts, these methods depended heavily on 3D annotations for part-contrastive learning, which confines them to a narrow set of categories and hinders efficient scaling. In this work, we present DINeMo, a novel neural mesh model that is trained with no 3D annotations by leveraging pseudo-correspondence obtained from large visual foundation models. We adopt a bidirectional pseudo-correspondence generation method, which produce pseudo correspondence utilizing both local appearance features and global context information. Experimental results on car datasets demonstrate that our DINeMo outperforms previous zero- and few-shot 3D pose estimation by a wide margin, narrowing the gap with fully-supervised methods by 67.3%. Our DINeMo also scales effectively and efficiently when incorporating more unlabeled images, which demonstrate the advantages over supervised learning methods that rely on 3D annotations. Our project page is available here.
@@ 1 Introduction
In this work, we present DINeMo, a novel neural mesh model that is trained without 3D annotations , enabled by leveraging large pretrained visual foundation models, such as DINOv2 [20]. Rather than relying on groundtruth keypoint locations as in previous works [13, 9], we train DINeMo with part-contrastive loss using pseudo-correspondence obtained by matching neural features from SD-DINO [35]. However, raw pseudo-correspondence from SD-DINO can be quite noisy, e.g., keypoints are often mismatched between left and right. We argue that keypoint correspondence matching should consider both local information, i.e., per patch feature similarities, and global context information, i.e., 3D orientation of the object. Based on this motivation, we propose a novel bidirectional pseudo-correspondence generation, which consists of two steps: (i) matching a global pose label from raw keypoint correspondences, and (ii) refine local keypoint correspondences based on the predicted global pose label. Lastly we extend standard analysis-by-synthesis inference with Grounded-SAM[23] masks, achieving enhanced occlusion robustness.
Extensive results on car class demonstrate that our DINeMo outperforms previous zero- and few-shot 3D pose estimation methods by a wide margin on both in-distribution and partial occlusion testing data, achieving a 67.3% reduction in the performance gap relative to fully supervised methods. Our approach also outperforms all previous methods on SPair71k [16] for semantic correspondence.
As DINeMo can be trained on object images without 3D annotations, we further study the scaling properties of our approach by involving abundant data from public image datasets, such as Stanford Cars [12]. Experimental results demonstrate that our DINeMo scales effectively and efficiently by involving more unlabeld images for training. This demonstrates the advantages of DINeMo over previous fully-supervised approaches [13, 9], which struggle to scale up due to the reliance on 3D annotations.
In summary, our main contributions are as follows: (1) We present DINeMo, a novel neural mesh model trained with pseudo-labels from our bidirectional pseudo-correspondence generation method. (2) Experimental results show that DINeMo outperforms previous zero- and few-shot methods by a wide margin, largely narrowing the gap with fully-supervised methods. (3) Our DINeMo scales effectively with more unlabeled images for training.
@@ 2 Related Work
Multiple strategies have been developed for category-level 3D pose estimation. Traditional methods formulate this task as a classification problem that predicts discrete pose labels [28, 17]. Another line of work follows a two-stage, keypoint-driven pipeline [39], where semantic keypoints are detected and used to estimate pose via Perspective-n-Point algorithm. More recent advancements shifted toward render-and-compare paradigms [30, 13, 9, 33], transitioning from generating pixel values to semantic feature representations. These methods frame pose estimation as an optimization problem that minimizes the discrepancy between features extracted from input image and ones rendered from a posed 3D mesh.
Keypoint correspondence.
Several approaches have been developed to establish correspondences between 2D image keypoints and 3D mesh vertices—a fundamental component in tasks such as shape reconstruction and 3D pose estimation. At the category level, however, most existing methods focused on deformable object classes, particularly humans [4, 22] and animals [26]. This focus is largely attributed to the availability of large-scale datasets with dense image-to-template annotations, such as DensePose-COCO [4] for humans and DensePose-LVIS [19] for animals. Notably, recent work has demonstrated the feasibility of learning dense correspondences without explicit keypoint supervision [26], particularly in the animal domain. In this work, we extend this line of research to rigid object categories, which have received comparatively less attention.
Render-and-compare
approaches estimate 3D poses by minimizing the reconstruction error between feature representations projected from a 3D object and those extracted from the input image. This strategy can be interpreted as a form of approximate analysis-by-synthesis [3], which contrasts with purely discriminative methods and has demonstrated increased robustness to out-of-distribution scenarios. Particularly, such approaches have shown effectiveness in handling partial occlusions, proving beneficial in both object classification [11] and 3D pose estimation tasks [30, 8].
@@ 3 Methods
In this section we introduce our DINeMo, as illustrated in Fig. 1. We start by reviewing the neural mesh models in Sec. 3.1. Then we introduce our bidirectional pseudo-correspondence generation in Sec. 3.2. Lastly we present our inference method in Sec. 3.3, which leverages pretrained segmentation model to optimize pose parameters.
@@ 3.1 Preliminaries: Neural Mesh Models
where $f_{\Phi}$ is the network backbone parameterized by $\Phi$, $\mathcal{FG}$ and $\mathcal{BG}$ are set of foreground and background positions, and $C_{b}$ is the background feature. The network parameters $\Phi$ and learnable feature $\{\mathcal{C},C_{b}\}$ are optimized with a part-contrastive loss. During inference, we minimize the negative log-likelihood w.r.t. pose parameters $m$.
@@ 3.2 Bidirectional Pseudo-Correspondence
Previous works train neural mesh models on keypoint correspondences obtained from 3D annotations, such as object or human poses. In this work, we propose to leverage pseudo-correspondence from visual foundation models, eliminating the need for 3D annotations that are difficult to obtain.
@@ 3.3 Inference
To find foreground and background regions, previous works [13, 9, 30] introduce an one-hot map $\mathcal{Z}$ indicating if each 2D location is visible or not based on feature activations. Empirically we find that the one-hot map often involves a significant portion of the background, leading to degraded performance on partial occlusion data. Given the recent advancements of segmentation methods, we first predict $\mathcal{Z}$ with a SAM2 [21] model and then optimizes the negative log-likelihood loss while fixing $\mathcal{Z}$.
@@ 4.2 Scaling Properties
As our DINeMo is trained on object images without 3D annotations, we study the scaling properties w.r.t. different training data sizes. Specifically, we trained a variety of DINeMo models on different numbers of unlabeled images, ranging from 2048 images (comparable to the training set size in PASCAL3D+ [32]) up to 15,000 images (as in Stanford Cars dataset [12]). As shown from the results in Fig. 4, DINeMo scales well with more unlabeled images used during training, i.e., the pose accuracy at $\pi/6$ increases from 93.1% to 93.7% and the per point PCK increases from 62.3 to 64.2. This highlights the advantages over previous supervised learning methods — DINeMo does not require 3D annotations that are hard to obtain, and scales effectively and efficiently by involving more unlabeled images for training, which are abundant on the Internet [25].
@@ 5 Conclusions
In this work we present DINeMo, a novel neural mesh model that is trained with no 3D annotations-only from pseudo-correspondence obtained from large visual foundation models. We propose a novel bidirectional pseudo-correspondence generation method that can effectively utilize both local appearance features and global context information to produce more 3D-consistent pseudo-correspondence. Experimental results on car datasets demonstrate that our DINeMo outperforms previous zero- and few-shot 3D pose estimation methods by a wide margin, narrowing the gap with fully-supervised methods by 67.3%. By incorporating more unlabeled images during training, our DINeMo also scales effectively and efficiently, demonstrating the advantages over fully supervised methods that rely on scarce 3D annotations.
6 Acknowledgements
AY acknowledges support via Army Research Laboratory award W911NF2320008 and Office of Naval Research with award N000142412696.